# SeFi-IDE: Semantic-Fidelity Identity Embedding for Personalized Diffusion-Based Generation

### 31 Jan 2024 v1

@@ 4.4 Generalization: Multi-ID and Other Objects
We further validate our methods on generation of multiple IDs. The results in Fig. 8 show that training individually can also perform good performance of human-human interactions. Furthermore, compared to Celeb Basis [47], our method does not introduce face prior from other models (e.g., a pre-trained face recognition model or basis module). As shown in Fig. 9, we adopt bear, cat, and dog for experiments, which show the generalization ability of our method.
5 Conclusion
@@
We propose two novel problem-orient techniques to enhance the semantic-fidelity of ID embedding for personalized diffusion-based generation. We analyze the attention overfit problem and propose Face-Wise Attention Loss. This improves the ID accuracy and facilitates the effective interactions between this ID embedding and other concepts (e.g., scenes, facial attributes, and actions). Then, we optimize one ID embedding as multiple per-stage tokens, which further expands the textual conditioning space with semantic-fidelity control ability. Extensive experiments validate our better ID accuracy and manipulation ability than previous methods, and we thoroughly conduct ablation study to validate the effectiveness of our methods.

### 22 Mar 2024 v2

@@ IV-I Generalization
Embedding Other Objects. We further validate our methods on other objects. Compared to Celeb Basis [3], our method does not introduce face prior from other models (e.g., a pre-trained face recognition model or basis module). As shown in Fig. 15, we adopt animals (Bear, Cat, and Dog) and general objects (Car, Chair and Plushie) for experiments, which show the generalizability of our method.
Using Stable Diffusion XL. To validate the generalization to the latest version of Stable Diffusion Model, we select SDXL model [69] stable-diffusion-xl-base-1.0 as the target model and the newly released methods using it for comparisons. As shown in Fig. 16, we compare with the SOTA methods InstantID [57], PhotoMaker [59], and IP-Adapter-FaceID [68]. InstantID [57] can only generate the face photo and fails to manipulate other actions or facial attributes. Although PhotoMaker [59] and IP-Adapter-FaceID [68] could generate the target ID under different scenes and actions, they can not handle complex actions (e.g., “sit on a chair”) and accurate facial attribute controlling. Additionally, IP-Adapter-FaceID [68] even loses the identity information of target person when combined with facial attribute prompts. As shown in the shortcomings of other methods, we found that incorporating additional features into the SDXL model would compromise the semantic-fidelity ability of T2I models, resulting in generated images that are misaligned with the given prompts. In contrast, our approach focuses on learning interactive ID embeddings with diffusion prior itself, which would not disrupt the original semantic understanding capability of the adopted models.
V Conclusion
@@
We propose two novel problem-orient techniques to enhance the accuracy and interactivity of the ID embeddings for semantic-fidelity personalized diffusion-based generation. We analyze the attention overfit problem and propose Face-Wise Attention Loss. This improves the ID accuracy and facilitates the effective interactions between this ID embedding and other concepts (e.g., scenes, facial attributes, and actions). Then, we optimize one ID embedding as multiple per-stage tokens, which further expands the textual conditioning space with semantic-fidelity control ability. Extensive experiments validate our better ID accuracy and manipulation ability than previous methods, and we thoroughly conduct ablation study to validate the effectiveness of our methods. Moreover, our embedding method does not rely on any prior facial knowledge, which is potential to be applied to other categories.
Ethical Statement. Our research endeavors are dedicated to addressing specific challenges within multi-modal generation with the overarching aim of advancing the technological landscape within our community. We staunchly oppose any misuse of our technology, such as the unauthorized use of their identity information. To mitigate such risks, we are actively developing watermarking techniques to prevent the misuse of Artificial Intelligence Generated Content.